OpenStudy (anonymous):

You put $5,000 in a simple interest account at a bank. You will earn $1,200 in just 4 years. What is the annual interest rate for the account?

OpenStudy (anonymous):

As given Principal,the amount u deposited =$5000,Interest u earned,I=$1200 and time it took,n=4 yrs So using simple interest formula, I=P*r*n/100; where r is the rate so, put the values 1200=5000*r*4/100 1200*100=5000*r*4 r=(1200*100)/(5000*4) r=6% p.a.